OpenStudy (my4babiesandme):

find the following assume variable can be repersented by any real numbers. Sqrt (a+9)^2

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{(a+9)^2}=|a+9|\]

OpenStudy (anonymous):

the sqrt and the ^2 would cancel out, so it would be the absolute value of a + 9

OpenStudy (my4babiesandme):

thank you

jimthompson5910 (jim_thompson5910):

This is using the idea that \[\large \sqrt{x^2}=|x|\]